The base of a triangle is 2 feet more than twice its altitude. If the area is 12 square feet, find the dimensions.
Answer by checkley77(12844)

B=2H+2
A=BH/2
12=(2H+2)H/2
12=(2H^2+2H)/2
2H^2+2H=2*12
2H^2+2H-24=0
2(H^2+H-12)=0
2(H+4)(H-3)=0
H-3=0
H=3 ANS. FOR THE HEIGHT.
B=2*3+2
B=6+2=8 ANS. FOR THE BAS.
PROOF:
12=8*3/2
12=24/2
12=12